# Compactly supported $\mathbb{A}^{1}$-Euler characteristics of symmetric powers of cellular varieties

### v1

@@ Abstract.
The compactly supported $\mathbb{A}^{1}$-Euler characteristic, introduced by Hoyois and later refined by Levine and others, is an anologue in motivic homotopy theory of the classical Euler characteristic of complex topological manifolds. It is an invariant on the Grothendieck ring of varieties $\mathrm{K}_{0}(\mathrm{Var}_{k})$ taking values in the Grothendieck-Witt ring $\mathrm{GW}(k)$ of the base field $k$. The former ring has a natural power structure induced by symmetric powers of varieties. In a recent preprint, Pajwani and Pál construct a power structure on $\mathrm{GW}(k)$ and show that the compactly supported $\mathbb{A}^{1}$-Euler characteristic respects these two power structures for 00-dimensional varieties, or equivalently étale $k$-algebras. In this paper, we define the class $\mathrm{Sym}_{k}$ of symmetrisable varieties to be those varieties for which the compactly supported $\mathbb{A}^{1}$-Euler characteristic respects the power structures and study the algebraic properties of $\mathrm{K}_{0}(\mathrm{Sym}_{k})$. We show that it includes all cellular varieties, and even linear varieties as introduced by Totaro. Moreover, we show that it includes non-linear varieties such as elliptic curves. As an application of our main result, we compute the compactly supported $\mathbb{A}^{1}$-Euler characteristics of symmetric powers of Grassmannians and certain del Pezzo surfaces.
@@ 1. Introduction
The compactly supported $\mathbb{A}^{1}$-Euler characteristic $\chi^{\mathrm{mot}}$ was first introduced in work of Hoyois [Hoy14], later refined by Levine [Lev20] for smooth projective schemes and extended to general varieties over a field in characteristic zero by Arcila-Maya, Bethea, Opie, Wickelgren and Zakharevich [AMBO+22], and to general varieties in characteristic not equal to $2$ by Levine, Pepin-Lehalleur and Srinivas in [LPS24]. It is an algebro-geometric invariant that refines both the real and complex Euler characteristic of topological manifolds, as well as some additional arithmetic data. As opposed to the classical Euler characteristic, which takes values in $\mathbb{Z}$, the compactly supported $\mathbb{A}^{1}$-Euler characteristic takes values in the Grothendieck-Witt ring $\operatorname{GW}(k)$ of the base field $k$, so it contains “quadratic” information. However, unlike the classical Euler characteristic, it can be difficult to compute $\chi^{\mathrm{mot}}(X)$ even when $X$ is a smooth projective variety. Papers such as [LPS24] and [Vie23] use the motivic Gauss-Bonnet Theorem of Levine-Raksit [LR20] to compute the compactly supported $\mathbb{A}^{1}$-Euler characteristic of hypersurfaces in $\mathbb{P}^{n}$ and complete intersections of hypersurfaces of the same degree in $\mathbb{P}^{n}$. Also, Brazelton, McKean and Pauli computed the compactly supported $\mathbb{A}^{1}$-Euler characteristics of Grassmannians in [BMP23], using $\mathbb{A}^{1}$-degrees. While this invariant can be difficult to work with, it has found use in enumerative geometry since it is analogous to the classical Euler characteristic of a manifold. We may use this invariant to obtain enumerative geometry counts which take values in $\operatorname{GW}(k)$, and papers such as [PP22] by Pajwani and Pál and [BBG24] by Blomme, Brugallé and Garay, use the compactly supported $\mathbb{A}^{1}$-Euler characteristic to obtain arithmetic refinements of results in complex enumerative geometry, the first over a general base field and the second over the real numbers.
@@
This paper is concerned with the compactly supported $\mathbb{A}^{1}$-Euler characteristic of symmetric powers of varieties. These geometric objects are closely related to Hilbert schemes of points via the birational Hilbert-Chow morphism. They are of particular interest to people studying enumerative geometry, appearing for example in the Göttsche formula for Euler characteristics of Hilbert schemes of surfaces ([Göt90, Theorem 0.1], [PP22, Corollary 8.5]). Since these varieties are almost always singular if $\mathrm{dim}(X)\geq 2$, we cannot directly apply the motivic Gauss–Bonnet theorem of [LR20] to them, and as such their compactly supported $\mathbb{A}^{1}$-Euler characteristics seem difficult to compute directly. Therefore, we instead use the power structures defined in [PP23] for this computation. We give a formula for the compactly supported $\mathbb{A}^{1}$-Euler characteristic of symmetric powers of a large class of varieties that we call $\operatorname{K}_{0}$-étale linear, see Definition 3.1. Informally, $\operatorname{K}_{0}$-étale linear varieties are varieties whose class in $\operatorname{K}_{0}(\mathrm{Var}_{k})$ decomposes into a sum with terms $[\mathbb{A}^{n}_{L}]$, where $L/k$ is a finite separable extension (see Definition 3.1). These form a large class of varieties containing many widely studied varieties, such as cellular varieties (Lemma 3.3), del Pezzo surfaces of degree $\geq 5$ (Theorem 5.7), certain tori (Example 3.2) and others. Our main result can be stated as follows:
@@
Let $X$ be a $\operatorname{K}_{0}$-étale linear variety over a base field $k$ of characteristic 00 (see Definition 3.1), and for $n\in\mathbb{Z}_{\geq 0}$, write $X^{(n)}:=\mathrm{Sym}^{n}(X)$. Then $\chi^{\mathrm{mot}}(X^{(n)})=a_{n}(\chi^{\mathrm{mot}}(X))$ for every $n$, where $a_{n}$ denotes the function defining the power structure on $\operatorname{GW}(k)$ as in Definition 2.8.
The power of the above theorem lies in the fact that it is much easier to work with the power structure on $\operatorname{GW}(k)$ than it is to decompose the symmetric powers of $\operatorname{K}_{0}$-étale linear varieties in general.
We define a variety $X$ to be symmetrisable if $\chi^{\mathrm{mot}}$ respects the power structure as in our main result, i.e. $\chi^{\mathrm{mot}}(X^{(n)})=a_{n}(\chi^{\mathrm{mot}}(X))$ for all $n$, see Definition 4.1. Corollaries 6.4 and 6.6 show that the class of symmetrisable varieties contains curves of genus $1$ and that these are not $\operatorname{K}_{0}$-étale linear.
Theorem 1.2 (Corollary 6.4 and Corollary 6.6).
Let $C$ be a curve of genus $1$. Then $\chi^{\mathrm{mot}}(C^{(n)})=a_{n}(\chi^{\mathrm{mot}}(C))$, but $[C]\not\in\operatorname{K}_{0}(\mathrm{\acute{E}tLin}_{k})$
Additionally, we show in Theorem 4.17 that a variety over $k$ must itself be symmetrisable if it becomes symmetrisable after base change to a finite extension $L/k$ of odd degree. We use this to show that even dimensional Severi–Brauer varieties are symmetrisable in Corollary 4.19 even though they may not be $\operatorname{K}_{0}$-étale linear. While we only show that curves of genus $\leq 1$ are symmetrisable, in forthcoming work, Lukas Bröring and the third author [BV24] show that all curves are symmetrisable using different techniques.
We apply our main result in Theorem 5.8 to compute $\chi^{\mathrm{mot}}(X^{(3)})$ for $X$ a cubic surface; a computation which we believe would be difficult to do without using the power structure. Similarly, we use it to compute a generating series for $\chi^{\mathrm{mot}}$ of the symmetric powers of a Grassmannian.
@@
$\sum_{t=0}^{\infty}\chi^{mot}(\mathrm{Gr}(d,r)^{(n)})t^{n}=(1-t)^{-e(d,r)}(1-(\langle-1\rangle t))^{-o(d,r)}\in\operatorname{GW}(k)[[t]],$
where $e(d,r)$ is the $d$-th entry in the $r$-th row of Losanitsch’s triangle, and $o(d,r)=\binom{r}{d}-e(d,r)$.
The above result enriches the generating series of the classical Euler characteristic of symmetric powers of Grassmannians, as the rank map $\operatorname{GW}(k)\rightarrow\mathbb{Z}$ sends the form $\langle-1\rangle$ to $1$ and the sum $e(d,r)+o(d,r)$ is the binomial coefficient $\binom{r}{d}$.
In Section 2, we recall notions required for our paper. We first restate the definition of the compactly supported $\mathbb{A}^{1}$-Euler characteristic in Definition 2.3. To compute the compactly supported $\mathbb{A}^{1}$-Euler characteristics of symmetric powers of varieties, we use the notion of a power structure on a ring, see Definition 2.6. We recall the existence natural power structures on both $\operatorname{K}_{0}(\mathrm{Var}_{k})$ and $\operatorname{GW}(k)$ following [GZLMH06] and [PP23]. We introduce the notion of a $\operatorname{K}_{0}$-étale linear variety in Section 3 (Definition 3.1), and prove some of their basic properties. Section 4 is concerned with proving the main theorem of this paper, using Göttsche’s lemma for symmetric powers [Göt01, Lemma 4.4]. Section 5 then uses the main result to compute the Euler characteristics of Grassmannians and a sizeable class of del Pezzo surfaces. Finally in Section 6, we turn our attention to varieties which do not become $\operatorname{K}_{0}$-étale linear over any field, but are nonetheless symmetrisable.
@@ Acknowledgements
The first author was funded by the Marsden grant “Rational points and Anabelian Geometry”. The second author was supported by the ERC through the project QUADAG. This paper is part of a project that has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 832833).
The authors thank Marc Levine for pointing out the elegant proof of Proposition 6.3 and some errors in an earlier version of this preprint, as well as H. Uppal for useful conversations about del Pezzo surfaces. We also thank Margaret Bilu for helpful disucssions on the paper [BHS+23], and Ambrus Pál for helpful discussions and worked examples which agree with the main theorem. We thank Lukas Bröring for helpful discussions on the positive characteristic case and for many useful comments on an earlier draft of this paper. We further thank Stefan Schreieder for proofreading an earlier draft of this paper.
Notation
Fix $k$ to be a field of characteristic 00. For a variety $X$ over $k$, i.e. a reduced separated scheme of finite type over $k$, and $n\in\mathbb{Z}_{\geq 0}$, let $X^{(n)}$ be the $n^{\text{th}}$ symmetric power of $X$, which is the quotient of $X^{n}$ by the action of the symmetric group on $n$ letters permuting the co-ordinates.
@@ 2.1. The compactly supported $\mathbb{A}^{1}$-Euler characteristic
Let $\mathrm{Var}_{k}$ be the category of varieties over $k$. The Grothendieck ring of varieties $\operatorname{K}_{0}(\mathrm{Var}_{k})$ is the free abelian group generated by isomorphism classes $[X]$ of varieties $X\in\mathrm{Var}_{k}$ modulo the relation $[X]=[Z]+[X\setminus Z]$ for every closed immersion $Z\rightarrow X$ in $\mathrm{Var}_{k}$, together with the multiplication given on generators by $[X][Y]=[X\times_{k}Y]$. Note that $1=[\operatorname{Spec}k]$ and $0=[\emptyset]$ in $\operatorname{K}_{0}(\mathrm{Var}_{k})$. Denote the subring of $\operatorname{K}_{0}(\mathrm{Var}_{k})$ which is generated by dimension 00 varieties by $\operatorname{K}_{0}(\mathrm{\acute{E}t}_{k})$.
@@
To define $\chi^{\mathrm{mot}}$, we follow [LR20, Corollary 8.7] and [AMBO+22, Definition 1.4, Theorem 1.13]. For $X$ a smooth projective scheme over $k$ of dimension $n$, define a quadratic form $\chi^{Hdg}(X)\in\operatorname{GW}(k)$ as follows.
  • If $n$ is odd, we set $\chi^{Hdg}(X)=m\cdot H$ where
@@
  • If $n=2p$ is even, we set $\chi^{Hdg}(X)=m\cdot H+Q$ where $Q$ corresponds to the symmetric bilinear form given by
[…]
@@
By [AMBO+22, Theorem 1.13], there exists a unique ring homomorphism, the compactly supported $\mathbb{A}^{1}$-Euler characteristic
$\chi^{\mathrm{mot}}_{k}:\operatorname{K}_{0}(\mathrm{Var}_{k})\to\operatorname{GW}(k),$
such that if $X$ is a smooth projective connected variety, $\chi^{\mathrm{mot}}_{k}([X])=\chi^{Hdg}(X)$.
@@
For a variety $X$ over $k$, the compactly supported $\mathbb{A}^{1}$-Euler characteristic $\chi^{\mathrm{mot}}_{k}(X)\in\operatorname{GW}(k)$ is the image of $[X]\in\operatorname{K}_{0}(\mathrm{Var}_{k})$ under the above map.
@@
When the base field is clear, we will drop the subscript $k$ and simply write $\chi^{\mathrm{mot}}(X)$ to mean $\chi^{\mathrm{mot}}_{k}([X])$.
In [AMBO+22], this invariant is denoted by $\chi^{\mathbb{A}^{1}}_{c}$.
@@
For $X$ smooth and projective, the motivic Gauss–Bonnet Theorem ([LR20, Theorem 1.3]) implies that $\chi^{\mathrm{mot}}(X)$ is the quadratic Euler characteristic of $X$. This is an invariant coming from motivic homotopy theory which was first studied by Hoyois in [Hoy14]. One obtains this invariant by applying the categorical Euler characteristic construction as defined by Dold-Puppe [DP80] to the stable motivic homotopy category $\operatorname{\mathbf{SH}}(k)$ introduced by Morel-Voevodsky, see [Lev20, Section 2] for details. The quadratic Euler characteristic above is the motivation for the definition of the compactly supported $\mathbb{A}^{1}$-Euler characteristic in [AMBO+22]. For this paper, we define $\chi^{\mathrm{mot}}_{k}$ in terms of Hodge cohomology for ease of use, however this invariant should be thought of as one coming from motivic homotopy theory.

### v2

@@ Abstract.
The compactly supported $\mathbb{A}^{1}$-Euler characteristic, introduced by Hoyois and later refined by Levine and others, is an analogue in motivic homotopy theory of the classical Euler characteristic of complex topological manifolds. It is an invariant on the Grothendieck ring of varieties $\mathrm{K}_{0}(\mathrm{Var}_{k})$ taking values in the Grothendieck-Witt ring $\mathrm{GW}(k)$ of the base field $k$. The former ring has a natural power structure induced by symmetric powers of varieties. In a recent preprint, the first author and Pál construct a power structure on $\mathrm{GW}(k)$ and show that the compactly supported $\mathbb{A}^{1}$-Euler characteristic respects these two power structures for 0-dimensional varieties, or equivalently étale $k$-algebras. In this paper, we define the class $\mathrm{Sym}_{k}$ of symmetrisable varieties to be those varieties for which the compactly supported $\mathbb{A}^{1}$-Euler characteristic respects the power structures and study the algebraic properties of the subring $\mathrm{K}_{0}(\mathrm{Sym}_{k})$ of symmetrisable varieties. We show that it includes all cellular varieties, and even linear varieties as introduced by Totaro. Moreover, we show that it includes non-linear varieties such as elliptic curves. As an application of our main result, we compute the compactly supported $\mathbb{A}^{1}$-Euler characteristics of symmetric powers of Grassmannians and certain del Pezzo surfaces.
Key words and phrases:
motivic homotopy theory, refined enumerative geometry, Euler characteristics, symmetric powers, cellular varieties
@@ 1. Introduction
The compactly supported $\mathbb{A}^{1}$-Euler characteristic $\chi_{c}^{\mathrm{mot}}$ was first introduced in work of Hoyois [Hoy14], later refined by Levine [Lev20] for smooth projective schemes and extended to general varieties over a field in characteristic zero by Arcila-Maya, Bethea, Opie, Wickelgren and Zakharevich [AMBO+22] and Röndigs [Rön25], and to general varieties in characteristic not $2$ by Levine, Pepin-Lehalleur and Srinivas in [LPS24]. It is an algebro-geometric invariant that refines both the real and complex Euler characteristic of topological manifolds, as well as some additional arithmetic data. As opposed to the classical Euler characteristic, which takes values in $\mathbb{Z}$, the compactly supported $\mathbb{A}^{1}$-Euler characteristic takes values in the Grothendieck-Witt ring $\operatorname{GW}(k)$ of the base field $k$, so it contains “quadratic” information. However, unlike the classical Euler characteristic, it can be difficult to compute $\chi_{c}^{\mathrm{mot}}(X)$ even when $X$ is a smooth projective variety. Papers such as [LPS24] and [Vie25] use the motivic Gauß-Bonnet Theorem of Levine-Raksit [LR20] to compute the compactly supported $\mathbb{A}^{1}$-Euler characteristic of hypersurfaces in $\mathbb{P}^{n}$ and complete intersections of hypersurfaces of the same degree in $\mathbb{P}^{n}$, and Brazelton, McKean and Pauli computed the compactly supported $\mathbb{A}^{1}$-Euler characteristics of Grassmannians in [BMP23], using $\mathbb{A}^{1}$-degrees. While this invariant can be difficult to work with, it has found use in enumerative geometry since it is analogous to the classical Euler characteristic of a manifold. We may use this invariant to obtain enumerative geometry counts which take values in $\operatorname{GW}(k)$ and papers such as [PPar] by the first author and Pál, and [BBGar] by Blomme, Brugallé and Garay, use the compactly supported $\mathbb{A}^{1}$-Euler characteristic to obtain arithmetic refinements of results in complex enumerative geometry, the first over a general base field and the second over the real numbers.
@@
This paper is concerned with the compactly supported $\mathbb{A}^{1}$-Euler characteristic of symmetric powers of varieties, which can be viewed as moduli spaces of effective zero-cycles. These geometric objects are closely related to Hilbert schemes of points via the birational Hilbert-Chow morphism. They are of particular interest to people studying enumerative geometry, appearing for example in the Göttsche formula for Euler characteristics of Hilbert schemes of surfaces ([Göt90, Theorem 0.1], [PPar, Corollary 8.18]). Since symmetric powers of a variety $X$ are almost always singular if $\mathrm{dim}(X)\geq 2$, we cannot directly apply the motivic Gauß–Bonnet theorem of [LR20] to them, and as such their compactly supported $\mathbb{A}^{1}$-Euler characteristics seem difficult to compute directly. However, symmetric powers of a variety furnish an additional structure on $\mathrm{K}_{0}(\mathrm{Var}_{k})$, known as a power structure (see Definition 2.9). Therefore, we instead use the results of [PP25] to utilise a power structures defined on $\mathrm{GW}(k)$ in order to compute the motivic Euler characteristics of symmetric powers. We give a formula for the compactly supported $\mathbb{A}^{1}$-Euler characteristic of symmetric powers of a class of varieties that we call $\mathrm{K}_{0}$-étale linear. Informally, $\mathrm{K}_{0}$-étale linear varieties are varieties whose class in $\mathrm{K}_{0}(\mathrm{Var}_{k})$ decomposes into a sum with terms $\mathbb{A}^{n}_{L}]$, where $L/k$ is a finite separable extension (see Definition 3.1). These form a class of varieties containing many widely studied varieties, such as cellular varieties (Lemma 3.5), del Pezzo surfaces of degree $\geq 5$ (Theorem 5.7), certain tori (Example 3.2) and others. Our main result can be stated as follows:
@@
Let $X$ be a $\mathrm{K}_{0}$-étale linear variety over field $k$ of characteristic $\neq 2$ (see Definition 3.1), and for $n\in\mathbb{Z}_{\geq 0}$, write $X^{(n)}:=\mathrm{Sym}^{n}(X)$. Then $\chi_{c}^{\mathrm{mot}}(X^{(n)})=a_{n}(\chi_{c}^{\mathrm{mot}}(X))$ for every $n$, where $a_{n}$ denotes the function defining the power structure on $\operatorname{GW}(k)$ as in Definition 2.11.
The power of the above theorem lies in the fact that it is much easier to work with the power structure on $\operatorname{GW}(k)$ than it is to decompose the symmetric powers of $\mathrm{K}_{0}$-étale linear varieties in general.
We say a variety $X$ is symmetrisable if $\chi_{c}^{\mathrm{mot}}$ respects the power structure as in our main result, i.e., if $\chi_{c}^{\mathrm{mot}}(X^{(n)})=a_{n}(\chi_{c}^{\mathrm{mot}}(X))$ for all $n$, see Definition 4.1. Corollary 6.5 and Lemma 6.6 show that the class of symmetrisable varieties contains curves of genus $1$ and that these are not $\mathrm{K}_{0}$-étale linear.
Theorem 1.2 (Corollary 6.5 and Lemma 6.6).
Let $C$ be a curve of genus $1$, and let $k$ be a field of characteristic 0. Then $C$ is symmetrisable, but it is not $\mathrm{K}_{0}$-étale linear.
Additionally, we show in Theorem 4.16 that a variety over $k$ must itself be symmetrisable if it becomes symmetrisable after base change to a finite extension $L/k$ of odd degree. We use this to show that even dimensional Severi–Brauer varieties are symmetrisable in Corollary 4.18 even though they may not be $\mathrm{K}_{0}$-étale linear. While we only show that curves of genus $\leq 1$ are symmetrisable, in [BV25, Proposition 26], Bröring and the third author show that all curves are symmetrisable using different techniques. Moreover, in [Brö26, Theorem 8.3, Theorem 8.6], it is shown that for any smooth projective variety, $\chi_{c}^{\mathrm{mot}}(X^{(n)})=a_{n}(\chi_{c}^{\mathrm{mot}}(X))$ whenever $n\leq 3$, and the smooth projective assumption can be removed in characteristic 0 (see [Brö26, Theorem 8.9]).
We apply our main result in Theorem 5.9 to compute $\chi_{c}^{\mathrm{mot}}(X^{(3)})$ for $X$ a specific cubic surface; a computation which we believe would be difficult to do without using the power structure. Similarly, we use it to compute a generating series for $\chi_{c}^{\mathrm{mot}}$ of the symmetric powers of a Grassmannian.
@@
$\sum_{n=0}^{\infty}\chi_{c}^{\mathrm{mot}}(\mathrm{Gr}(d,r)^{(n)})t^{n}=(1-t)^{-e(d,r)}(1-(\langle-1\rangle t))^{-o(d,r)}\in\operatorname{GW}(k)[[t]],$
where $e(d,r)$ is the $d$-th entry in the $r$-th row of Losanitsch’s triangle, and $o(d,r)$ is given by $\binom{r}{d}-e(d,r)$.
The above result enriches the generating series of the classical Euler characteristic of symmetric powers of complex Grassmannians, as the rank map $\operatorname{GW}(k)\rightarrow\mathbb{Z}$ sends the form $\langle-1\rangle$ to $1$ and the sum $e(d,r)+o(d,r)$ is the binomial coefficient $\binom{r}{d}$. Applying the sign map $\langle-1\rangle\mapsto-1$ to the formula of Theorem 5.1 yields 0 if $r$ is even and $d$ is odd and $\binom{\lfloor r/2\rfloor}{\lfloor d/2\rfloor}$ otherwise, which is precisely the classical Euler characteristic of the real Grassmannian $\operatorname{Gr}(d,r)/\mathbb{R}$, in concordance with [Lev20, Remark 2.3.1].
In Section 2, we recall notions required for our paper. We first restate the definition of the compactly supported $\mathbb{A}^{1}$-Euler characteristic in Definition 2.5. To compute the compactly supported $\mathbb{A}^{1}$-Euler characteristics of symmetric powers of varieties, we use the notion of a power structure on a ring, see Definition 2.9. We recall the existence of natural power structures on both $\mathrm{K}_{0}(\mathrm{Var}_{k})$ and $\operatorname{GW}(k)$ following [GZLMH06] and [PP25]. We introduce the notion of a $\mathrm{K}_{0}$-étale linear variety in Section 3 (Definition 3.1), and prove some of their basic properties. Section 4 is concerned with proving the main theorem of this paper, using Göttsche’s lemma for symmetric powers [Göt01, Lemma 4.4]. Section 5 then uses the main result to compute the Euler characteristics of Grassmannians and a sizeable class of del Pezzo surfaces. Finally in Section 6, we turn our attention to varieties which do not become $\mathrm{K}_{0}$-étale linear over any field, but are nonetheless symmetrisable.
Notation
Fix $k$ to be a field of characteristic $\neq 2$. For a variety $X$ over $k$, i.e. a reduced separated scheme of finite type over $k$, and $n\in\mathbb{Z}_{\geq 0}$, let $X^{(n)}$ be the $n^{\text{th}}$ symmetric power of $X$, which is the quotient of $X^{n}$ by the action of the symmetric group on $n$ letters permuting the coordinates.
@@ Acknowledgements
This work was undertaken while the first author was funded by the Marsden grant “Rational points and Anabelian Geometry”, and he wishes to thank the University of Canterbury, New Zealand. The second author was supported by the ERC through the project QUADAG. This paper is part of a project that has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 832833).
The authors thank Marc Levine for pointing out the elegant proof of Proposition 6.4 and some errors in an earlier version of this preprint, as well as H. Uppal for useful conversations about del Pezzo surfaces. We also thank Margaret Bilu for helpful disucssions on the paper [BHS+24], and Ambrus Pál for helpful discussions and worked examples which agree with the main theorem. We thank Louisa Bröring for helpful discussions on the positive characteristic case and for many useful comments on an earlier draft of this paper. We further thank Stefan Schreieder for proofreading an earlier draft of this paper.
@@ 2.1. The compactly supported $\mathbb{A}^{1}$-Euler characteristic
Let $\mathrm{Var}_{k}$ be the category of varieties over $k$. The Grothendieck ring of varieties $\mathrm{K}_{0}(\mathrm{Var}_{k})$ is the free abelian group generated by isomorphism classes $[X]$ of varieties $X\in\mathrm{Var}_{k}$ modulo the relation $X]=[Z]+[X\setminus Z]$ whenever $Z\rightarrow X$ is a closed immersion, together with the multiplication given on generators by $X][Y]=[X\times_{k}Y]$. Note that $1=[\operatorname{Spec}k]$ and $0=[\emptyset]$ in $\mathrm{K}_{0}(\mathrm{Var}_{k})$. Denote the subring of $\mathrm{K}_{0}(\mathrm{Var}_{k})$ which is generated by dimension 0 varieties by $\mathrm{K}_{0}(\mathrm{\acute{E}t}_{k})$.
Following [CLNS18, Chapter 2, §4.4] and [BM25, §5], we also define a modified version of $\mathrm{K}_{0}(\mathrm{Var}_{k})$. Let $f:X\to Y$ be a morphism of varieties. We say that $f$ is a universal homeomorphism if for every $Y^{\prime}\to Y$, the induced map $f:X\times_{Y}Y^{\prime}\to Y^{\prime}$ is a homeomorphism of the underlying topological spaces of the schemes. Let $I^{uh}_{k}$ denote the ideal of $\mathrm{K}_{0}(\mathrm{Var}_{k})$ generated by classes of the form [X][Y][X]-[Y] for any pair of varieties $X,Y$ such that there exists a universal homeomorphism between $X$ and $Y$. Define $\mathrm{K}_{0}^{uh}(\mathrm{Var}_{k}):=\mathrm{K}_{0}^{uh}(\mathrm{Var}_{k})/I^{uh}_{k}$.
Remark 2.2.
As noted in [CLNS18, Chapter 2, Corollary 4.4.7], if $k$ has characteristic 0, then $I^{uh}_{k}=\{0\}$. In particular, in characteristic 0, $\mathrm{K}_{0}(\mathrm{Var}_{k})=\mathrm{K}_{0}^{uh}(\mathrm{Var}_{k})$. Results later in the paper often use a strategy of pushing down to $\mathrm{K}_{0}^{uh}(\mathrm{Var}_{k})$, proving the analogous result in this ring, and pulling back to $\mathrm{K}_{0}(\mathrm{Var}_{k})$, so in characteristic 0, these operations can be ignored.
@@
To define $\chi_{c}^{\mathrm{mot}}$, we follow [LR20, Corollary 8.7], [LPS24, Section 5.1] and [AMBO+22, Definition 1.4, Theorem 1.13]. For $X$ a smooth projective scheme over $k$ of dimension $n$, define a quadratic form $\chi^{\mathrm{Hdg}}(X)\in\operatorname{GW}(k)$ as follows:
  • If $n$ is odd, we set $\chi^{\mathrm{\mathrm{\mathrm{Hdg}}}}(X)=m\cdot H$ where
@@
  • If $n=2p$ is even, we set $\chi^{\mathrm{\mathrm{\mathrm{Hdg}}}}(X)=m\cdot H+Q$ where $Q$ corresponds to the non-degenerate symmetric bilinear form given by
    $H^{p}(X,\Omega^{p}_{X/k})\otimes H^{p}(X,\Omega^{p}_{X/k})\xrightarrow{\cup}H^{n}(X,\Omega^{n}_{X/k})\xrightarrow{\text{Trace}}k$
@@
Remark 2.4.
We note that the above quadratic form comes from the composition of cup product and trace (defined using Serre duality) which one can define on the Hodge cohomology groups $H^{i}(X,\Omega^{j}_{X/k})$ of $X$. Most of this quadratic form is hyperbolic; all of it if $n$ is odd, and everything except $Q$ if $n$ is even.
By [AMBO+22, Theorem 1.13] in characteristic zero and the discussion in [LPS24, Section 5.1] for more general fields, there exists a unique ring homomorphism, the compactly supported $\mathbb{A}^{1}$-Euler characteristic
$\chi^{\mathrm{mot}}_{c,k}:\mathrm{K}_{0}(\mathrm{Var}_{k})\to\operatorname{GW}(k),$
such that if $X$ is a smooth projective connected variety, $\chi_{c}^{\mathrm{mot}}([X])=\chi^{\mathrm{\mathrm{\mathrm{Hdg}}}}(X)$. Moreover, by [BM25, Corollary 5.4], $\chi^{mot}_{c,k}$ is 0 on $I^{uh}_{k}$, so $\chi^{mot}_{c,k}$ factors through $\mathrm{K}_{0}^{uh}(\mathrm{Var}_{k})$.
@@
For a variety $X$ over $k$, the compactly supported $\mathbb{A}^{1}$-Euler characteristic $\chi_{c}^{\mathrm{mot}}(X)\in\operatorname{GW}(k)$ is the image of $X]\in\mathrm{K}_{0}(\mathrm{Var}_{k})$ under the above map.
@@
When the base field is clear, we will drop the subscript $k$ and simply write $\chi_{c}^{\mathrm{mot}}(X)$ to mean $\chi^{\mathrm{mot}}_{c,k}([X])$. In [AMBO+22], this invariant is denoted by $\chi^{\mathbb{A}^{1}}_{c}$ and in [LPS24], it is denoted by $\chi_{c}$.
@@
The compactly supported $\mathbb{A}^{1}$-Euler characteristic carries a lot of information: if $k\subset\mathbb{R}$ and $X$ is a smooth projective variety over $k$, then by [Lev20, Remark 2.3.1] we have that the rank of $\chi_{c}^{\mathrm{mot}}(X)$ is equal to the topological Euler characteristic of $X(\mathbb{C})$ whereas the signature of $\chi_{c}^{\mathrm{mot}}(X)$ is equal to the topological Euler characteristic of $X(\mathbb{R})$.
Remark 2.8.
For $X$ smooth and projective, the motivic Gauß–Bonnet Theorem [LR20, Theorem 1.3] (which assumes for $k$ to be a perfect field, but also holds over a non-perfect base field by [Lev20, Remark 2.1(2)]) implies that $\chi_{c}^{\mathrm{mot}}(X)$ is the quadratic Euler characteristic of $X$. This is an invariant coming from motivic homotopy theory which was first studied by Hoyois in [Hoy14]. One obtains this invariant by applying the categorical Euler characteristic construction as defined by Dold-Puppe [DP80] to the stable motivic homotopy category $\operatorname{\mathbf{SH}}(k)$ introduced by Morel-Voevodsky, see [Lev20, Section 2] for details. The quadratic Euler characteristic above is the motivation for the definition of the compactly supported $\mathbb{A}^{1}$-Euler characteristic in [AMBO+22] and [LPS24]. For this paper, we define $\chi_{c}^{\mathrm{mot}}$ in terms of Hodge cohomology for ease of use, however this invariant should be thought of as one coming from motivic homotopy theory.